OpenStudy (anonymous):

(x+2)^2=49, then x+2=7 True or false?

OpenStudy (anonymous):

true

jimthompson5910 (jim_thompson5910):

since something like (-2)^2 = 4 and (2)^2 = 4, this means that if x^2 = 4, then x = 2 or x = -2

jimthompson5910 (jim_thompson5910):

so what this means is that if (x+2)^2=49, then x+2=7 or x+2 = -7

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

you're welcome